OpenStudy (anonymous):

THE TEMPERATURE T OF FOOD PUT IN THE FREEZER IS T=700/(t^2+4t+10) where t is the time n hours. FIND THE CHANGE OF T IN RESPECT TO t at each of the following times a)t=1 b)t=3 c)t=5

OpenStudy (anonymous):

take the derivative and plug in the numbers

OpenStudy (anonymous):

can you please tell me how??

OpenStudy (anonymous):

derivative is \[T'=-\frac{1400(t+2)}{(t^2+4t+10)^2}\]

OpenStudy (anonymous):

the derivative of \[\frac{1}{f(t)}\] is \[-\frac{f'(t)}{f^2(t)}\] so it is quick in this case

OpenStudy (anonymous):

i used \[f(t)=t^2+4t+10\] \[f'(t)=2t+4=2(t+2)\] and got the answer that way. not you just have to substitute in numbers for t

OpenStudy (anonymous):

can u do it for me plase i amlost???